### 06 Mar 2024 v1

@@ 1.2 Linguistic Framework
Our schema maintains that all prosodic phenomena may be analyzed as variations, either hierarchical or orthogonal, of a very small number of IU prototypes [7]. The variations include four basic layers: information structure, attitude, emotion, and 3-5 sub-categories of conversation action and/or discourse functions. See Figure 1 for an illustration.
@@ 1.3 Related Work
In the domain of computational prosody, improving speech synthesis has been a subject of significant research (e.g., [26]). As for automated analysis of prosody, most have been aimed at detecting single phenomena, such as unit boundaries (e.g., [27, 28]), prominence/saliency [28, 29, 30], or specific dialogue-acts (e.g., [31]). Crucially, none of the above tackle the multi-layered nature of many prosodic events.
@@ 2.1 Contextuality and Scope
The effect of the wider context on meaning includes multi-unit prosodic patterns. Simple examples are appositions, list patterns [48], and prosodic bi-partites (such as if-then constructions; see samples here). Paragraphs, narratives, and formal addresses encompass a larger scale, which often presents nested structures (e.g., a list of events within a narrative). Those form the syntax of prosodic patterns, compared to [49].
@@ 4.1.1 Manual annotation
Table 1: The annotated data. 1a. Number and fraction of prosodic prototypes; 1b. Number and fraction of main speaker vs. interviewees (n=82); 1c. Number and fraction of emphasis tags (= the number of words annotated).
@@ 4.2.2 Prediction
Algorithm 1 Pseudo code of the inference procedure. This method enables only prosodic labeling. Note that next label holds the predicted label of a multiclass-multilabel combination
@@ 4.2.3 Validation/Evaluation
As mentioned above (section 4.2.1), we trained for single recognition tasks in order to compare the performance on a single task vs. the triple one. In addition, we tested three different representation methods of prosodic labels: (1) ‘raw’, which refers to special ‘words’ that were generated for the process; (2) ‘compact’, which refers to twelve labels that stand for the twelve combinations of prosodic tags; and (3) ‘bits’, which is similar to ‘raw’, and represents each prosodic feature by a single token (see the supplementary material).
@@ 5 Results
Table 3 shows that the model generalizes well across datasets and genres. It was more successful when employed on the TAL data than on the Interviews data (which was excluded from the training material), and specifically so in regard to IU boundary recognition.
@@
Table 4: Ira Glass - the show host - vs other speakers (n=49 in test set), results for WHISPER-Small.

### 13 Mar 2024 v2

@@ 1.2 Linguistic Framework
Our schema maintains that all prosodic phenomena may be analyzed as variations, either hierarchical or orthogonal, of a very small number of IU prototypes [7]. The variations include four basic layers: information structure, attitude, emotion, and 3-5 sub-categories of conversation action and/or discourse functions (see Figure 1).
@@ 1.3 Related Work
In the domain of computational prosody, improving speech synthesis has been a subject of significant research (e.g., [26]). As for automated analyses of prosody, most have been aimed at detecting single phenomena, such as unit boundaries (e.g., [27, 28]), prominence/saliency [28, 29, 30], or specific dialogue-acts (e.g., [31]). Crucially, none of the above tackle the multi-layered nature of many prosodic events.
@@ 2.1 Contextuality and Scope
The effect of the wider context on meaning includes multi-unit prosodic patterns. Simple examples are appositions, list patterns [48], and prosodic bi-partites (such as if-then constructions; see samples here). Paragraphs, narratives, and formal addresses encompass a larger scale, which often presents nested structures (e.g., a list of events within a narrative). Those form the syntax of prosodic patterns (cf. [49]).
@@ 4.1.1 Manual annotation
Table 1: The annotated data. 1a. Number and fraction of main speaker data vs. interviewees (n=82); 1b. Number and fraction of prosodic prototypes; 1c. Number and fraction of emphasis tags (= the number of words annotated).
@@ 4.2.2 Prediction
Algorithm 1 Pseudo code of the inference procedure. This method enables prosodic labeling only. Note that next_label holds the predicted label of a multiclass-multilabel combination
@@ 4.2.3 Validation/Evaluation
As mentioned above (section 4.2.1), we trained for single recognition tasks in order to compare the performance on a single task vs. the triple one. In addition, we tested three different representation methods of prosodic labels: (1) ‘raw’, which refers to special ‘words’ that were generated for this process; (2) ‘compact’, which refers to twelve labels that stand for the twelve combinations of prosodic tags; and (3) ‘bits’, which is similar to ‘raw’, and represents each prosodic feature by a single token (see supplementary material).
@@ 5 Results
Table 3 shows that the model generalizes well across datasets and genres. It was more successful when employed on the TAL data than on the Interviews data (which were excluded from the training material), and specifically so in regard to IU boundary recognition.
@@
Table 4: Ira Glass - the show host - vs. other speakers (n=49 in test set), results for WHISPER-Small.